t acuiues only a Lalculator to find exact values of n for each of Llle Pollowing interest rates, and comment on how closely the Rule-of-70 holds. (a) r 0.01 (b) r:-0.05 (c)0.10 (d) r 0.25 (e) r= 1.00 10. Consider the production function, Q = 15L45Ks, where Q is output, L is labor input, and K represents capital input. Using natural logarithms, transform this exponential function into a linear function. Now assume that L = 10 and K= 5. What is the value of In (Q)? Remembering that exp(In(Q) Q, determine the value of Q. 1. You are given $10,000 as a graduation present, and you want to buy a new car in the near future. Therefore you invest your gift in a conservative money market account paying a continuously compounded 5% per annum. Using the formula X(tn) X)e and natural logarithms, solve for n to determine how low it will take you to saveQuestion #10
Answer #1

Taking natural log of production function

Q= 15 L^4/5K1/5

LnQ= ln(15) +4/5(ln(L)+1/5(ln(K)

Putting values of k & L in above equation

LnQ= 2.70+ 4/5*2.30+1/5*1.61

LnQ= 2..70+1.84+ 0.32.

LnQ= 4.86

From the result it has been observed that the growth rate of production will be 4.86 when labor is increased by 10 units and capital by 5 units